Q1. The velocity of a particle moving along x-axis is directly proportional to x.at time t=0 x=2 and at t=10 x=4.Determine the position at t=5.


Q2.A bird is flying with velocity 5i+6j wrt wind and wind is blowing along x-axis with velocity u.If bird starts moving from A and after sometime reaches point B with ab making an angle of 37 deg with x-axis.Find the value of u.


 


 


 


 


 

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear aparna,

v=ax where a is a constant

dx/dt=ax

dx/x=adt

On integrating we get

lnx=at+c

 

By putting the values given we can calculate c

 and hence x...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

v=kx

 dx/dt =kx

  integrating both sides

 logx = kt + c ........1

 at t=0 , x is 2

  log2=c .......2

 at t=10, x is 4

   log4=10k + c

   2log2=10k + log2

     k=log2/10...,..3

 putting values of constants in eq 1

 we get logx=log2(t/10 +1)

at t =5

           logx=(3/2)log2

             x=2^3/2 units